# Variability of MHD Instabilities in Benign Termination of High-Current Runaway Electron Beams in the JET and DIII-D Tokamaks

### v1

@@ 1 Introduction
Mitigation schemes for plasma disruptions will be essential for future high-current tokamaks such as ITER or commercial fusion power plants, as they can impose significant thermal and mechanical loads on tokamak components [Loarte_2011]. A promising solution to mitigate the resulting runaway electron (RE) beams is the benign termination scheme, in which hydrogen is injected into the RE beam and a magnetohydrodynamic (MHD) instability deconfines the REs, increases the RE deposition wetted area, and terminates the RE beam benignly [pazsoldan2019ppcf, pazsoldan2021nf]. After being first observed on DIII-D [pazsoldan2019ppcf], dedicated experiments were performed on the Joint European Torus (JET) tokamak [reux2021prl, reux2022ppcf], along with the successful elaboration of this technique on the DIII-D [liu2019nf, pazsoldan2021nf], TCV, and AUG tokamaks [sheikh2024ppcf]. In these experiments, it was observed that there are limits to the amount of injected material for which the injection recombines the background, or companion, plasma. While on DIII-D this recombination was observed as a key component in making the mitigation both successful and benign, the corresponding density measurements on JET are not available for a number of discharges due to compromised diagnostic data and insufficient spatial resolution of the lines of sight of the interferometry systems. In addition, the interpretation of the pressure gauge signals was found to be hampered in these discharges [Schoonheere_2024]. However, all available data suggest that, also on JET, a successful recombination of the background plasma is the key requirement for a benign termination [reux2024runaway].
@@
First publications on the effect of material injections into RE beams in JET by Reux et al. used data collected before 2022 and, among other things, investigated the termination and mitigation characteristics in relation to the secondary injection material employed [reux2021prl, reux2022ppcf]. Not all of these early experiments were aimed at achieving a benign termination as defined by Paz-Soldan et al [pazsoldan2019ppcf]. In general, it was found that high-$Z$ or mixed secondary injections predominantly led to unsuccessful RE terminations. For such terminations, particularly with impurity fractions above 5%, a rich phenomenology was observed due to the variety of kinetic processes involved. Paz-Soldan et al. revisited the JET dataset and compared it with similar data from DIII-D [pazsoldan2021nf]. That work placed greater emphasis on investigating the MHD and equilibrium properties, although discharges with various secondary injection mixtures were also presented. At the time, it was concluded that successful and safe benign terminations were more likely to occur in the absence of high-$Z$ material. This finding motivated further JET experiments focusing on clean hydrogenic injections. The present publication is restricted to the subset of JET and DIII-D discharges with hydrogenic injections and impurity fractions below 5%, which have a higher fraction of succesful, benign terminations.
@@
Furthermore, all the JET results published so far were limited to discharges with pre-disruptive plasma currents below $1.5$ MA (JET discharge numbers below $100,000$). Recent experiments on JET in 2022 and 2023 have reached higher pre-disruptive plasma and RE currents than on other machines, up to $I_{p}\approx 3$ MA. They have revealed challenges in achieving benign termination at high pre-disruptive current values of $I_{p}\geq 2.5$ MA. This is illustrated in Fig. 1(a), which shows the pre-disruptive plasma current $I_{p}$ on the x-axis and the RE current prior to termination on the y-axis, with only non-benign, red cases for $I_{p}\approx 3.0$ MA on JET (r.h.s.). All cases with $I_{p}>1.5$ MA stem from JET campaigns in 2022 and 2023 and have not yet been published in peer-reviewed literature, though a short description can be found in Ref. [reux2024runaway]. In general, understanding the distinguishing features of these high-$I_{p}$ experiments on JET is critical for reactor-relevant extrapolation. In particular, studying the dynamics of the terminating MHD activity is crucial, as a conclusive explanation for what differentiates the nature of the terminating MHD event between benign and non-benign cases is still lacking. In addition, the referenced publications on DIII-D and JET, in particular by Paz-Soldan et al. [pazsoldan2021nf], highlighted a phenomenological difference with benign termination on DIII-D at the lowest edge safety factors (reproduced on AUG and TCV [sheikh2024ppcf]), while on JET, benign terminations at higher edge safety factors were also observed. This poses an open question of how observations from JET compare to those on DIII-D, TCV, and AUG, and how to unify the observed phenomenology.
@@ 2 Scenario and Methodology
Both approaches of driving the beam unstable can trigger a large-scale MHD event and terminate the RE beam either without (benign) or with (non-benign) localized impact of the RE beam on the first wall or divertor structure. This is observed after the onset of the termination (3), characterized by a drop in plasma current and safety factor, together with a spike in magnetic activity. In this work, most of the data points, if not otherwise explicitly mentioned, are sampled just prior to the termination. Specifically, this refers to a manually selected time point corresponding to the last measurement before the magnetic signature rises, marking the onset of the terminating MHD event. On JET, the classification of benign and non-benign terminations is based on IR camera observations and the corresponding heat fluxes, as well as on the scheme developed by Loarte et al. [Loarte_2011], which tracks the conversion of magnetic energy into kinetic energy, a process that is more efficient in non-benign cases. Both is documented in Reux et al. [reux2022ppcf]. On DIII-D, due to the lack of comparable IR camera coverage, measured hard X-ray (HXR) spikes serve as a indicator distinguishing clean, benign terminations from non-benign terminations in which the RE beam strikes the wall, emitting HXR, as already applied in [pazsoldan2021nf].
@@
This investigation is based on fast magnetic measurements for MHD mode decomposition and equilibrium reconstruction. The primary diagnostics for the magnetic measurements are the Mirnov coils, which are small pickup coils placed around the tokamak vessel to measure time-varying magnetic fields. They operate on the principle of electromagnetic induction: a changing magnetic flux through the coil induces a voltage proportional to $\dot{B}$, the time derivative of the perturbed magnetic field. An example of such a measurement during the terminating MHD event of the JET discharge #102617 is shown in Panel (a) of Fig. 3 (solid line). Arrays of Mirnov coils distributed toroidally enable the mode decomposition of the measured perturbations in terms of the relative phases and amplitudes for different mode numbers $n$ through matrix operations [Munaretto_2021]. While DIII-D provides a large number of magnetic arrays, at JET only 5577 fast Mirnov coils were usable, depending on the plasma discharge. This restricts the analysis to low mode numbers $n=1,2$. Because of their fast response and sampling rates on the MHz scale, Mirnov coils are well suited for studying MHD instabilities, mode growth, and termination dynamics, where rapid changes in the magnetic field occur. Integration of the coil signals over the MHD event yields $\delta B$, providing direct information on the amplitude of the underlying magnetic perturbations, as indicated by the shaded region under the curve in Panel (a). By fitting a linear curve, dotted line in Panel (b), to the logarithmic $\dot{B}$ values (solid line), the growth rate $\gamma$ can be measured. Fitting $\gamma$ with this approach is valid only during the initial exponential growth phase, when nonlinear coupling and saturation are not yet occurring. In this regime, fitting $\dot{B}$ or $\delta B$ yields identical results, differing only by constant offsets in the linear fit. The same constant offsets are introduced when considering conducting and shielding wall structures in front of magnetic sensors and, therefore, do not affect the results presented here. It should be noted that the Mirnov coil arrays used in this work measure the perturbed poloidal magnetic field $\delta B_{p}$, which, for simplicity, is referred to as $\delta B$ throughout this work. Normalization to the distance between the measurement array and the magnetic flux surface is found not to significantly alter the trends observed in the following analysis, thus this correction is not applied.
@@ 3.1 Evolution of high-$I_{p}$ RE beams on JET
The first critical observation concerns the fundamental difference in the duration of the RE plateau phase between benign and non-benign cases. In typical benign scenarios, the RE plateau phase is long, controlled, and stable, with an average duration of $0.62$ s (standard deviation $0.66$ s). In contrast, in cases where recombination is unsuccessful and the outcome is non-benign, no stable plateau forms, with an average duration of $0.07$ s (standard deviation $0.03$ s) before the termination. This is likely due to the higher resistivity of the non-recombined plasma, which results in a situation where the spatial position of the RE beam is no longer effectively controlled. The subsequent MHD response, however, is assumed to be largely indifferent to the specific pathway by which the system becomes unstable. A more detailed investigation of how the RE trajectory influences key quantities, such as the current profile distribution, and, consequently, the MHD instability is left for future work and cannot be disentangled from present experiments.
@@
The next step of the experimental analysis concerns the properties of the plasma equilibrium. In Fig. 4, the trajectory of the reconstructed magnetic axis is plotted for the last 300 ms before termination in the $R$-$z$-space with the JET contour in black for comparison. For cases in which the RE plateau phase was shorter, sampling was performed from the onset of the RE plateau phase. The panel on the left shows the non-benign cases, while the panel on the right shows the benign cases. The color coding corresponds to the cross-section of the RE column. It is observed that the onset of termination in non-benign cases occurs at a smaller cross-section and in a less central position, whereas benign terminations typically begin from a relatively larger cross-section and a more central position. This suggests a different evolution of the RE beam in benign and non-benign cases, with non-benign beams being compressed to smaller cross-sections and, consequently, to lower edge safety factors than their benign counterparts. As already noted, the non-benign cases on JET are usually very short, suggesting that the higher resistivity of non-recombined plasmas likely affects their controllability. Under these circumstances, the short lifetime of the non-recombined plasmas could interact with the evolution of the current profiles: assuming identical current-profile evolution across shots, any systematic variation in duration would alter the current profile at the point of the secondary disruption. This, as already mentioned, is subject of a subsequent publication on this topic.
@@
As discussed above, the non-benign cases on JET are associated with high pre-disruptive plasma currents. The higher stored energy in these cases leads to strong RE currents. Combined with the observed smaller cross-sections, this results in higher RE current densities prior to termination. Consequently, a clear separation between benign and non-benign cases emerges in terms of their RE current density $j_{RE}$, as illustrated in Fig. 1(b) with the JET datapoints on the r.h.s. This behavior is not commonly observed across the previously discussed devices, and a similar analysis for the studied DIII-D dataset does not show such a separation, see l.h.s. of the same figure. This suggests that different mechanisms play a role in the formation or termination of the RE currents on the two machines. Interestingly, at the highest $I_{p}\approx 1.2$ MA on DIII-D, the only non-benign case does separate from the benign cases with thresholds similar to those in the JET dataset, but more data would be needed to draw firm conclusions. For JET, however, the higher RE current densities may interact with kinetic effects, as suggested by Hollmann and Hoppe, who noted that more concentrated (peaked) RE current density profiles may hinder effective recombination of the companion plasma or can even lead to re-ionization of an initially recombined background plasma [Hollmann2023, Hoppe2025].
@@ 3.2 Intermittent MHD events prior to the termination in JET
In Fig. 6(a), the measured perturbed $\delta B$ is shown for benign and non-benign cases that pass through or terminate at $q_{\text{edge}}=3$. The MHD amplitude in benign cases increases linearly with the RE current deconfined during the event. This suggests that there exists a threshold in $\delta B/B$ (with $B=B_{\theta}\sim I_{RE}$) for which the MHD mode is strong enough to trigger a catastrophic, terminating MHD event. Non-benign cases exhibit a clearly suppressed MHD amplitude, insufficient to cause a proper termination. At the moment, it remains an open question how the recombination state of the companion plasma influences the amplitude or saturation of these MHD events.
A possible explanation for the observed suppression is provided in Panel (b), which shows a tendency toward higher $l_{i}$ values in most of the non-benign cases, suggesting a more peaked RE current distribution. Variations in the current profiles could originate from the dependence of the RE current density distribution on the specific mechanisms by which the RE current forms during the primary disruption. JET, in particular, is characterized by strong RE avalanche generation, which could explain the different $j_{RE}$ profiles observed. This was discussed in detail by Smith et al. [Smith_2010] who predicted that the absolute pre-disruptive current can influence the RE current distribution on JET. Steepening of the RE current profiles, correlated with higher pre-disruptive $I_{p}$, was also observed on JET by Loarte et al. [Loarte_2011].

### v2

@@ 1 Introduction
Mitigation schemes for plasma disruptions will be essential for future high-current tokamaks such as ITER or commercial fusion power plants, as they can impose significant thermal and mechanical loads on tokamak components [Loarte_2011]. A promising solution to mitigate the resulting runaway electron (RE) beams is the benign termination scheme, in which hydrogen is injected into the RE beam and a magnetohydrodynamic (MHD) instability deconfines the REs, increases the RE deposition wetted area, and terminates the RE beam benignly. After being first observed on DIII-D [pazsoldan2019ppcf], dedicated experiments were performed on the Joint European Torus (JET) tokamak [reux2021prl, reux2022ppcf], along with the successful elaboration of this technique on DIII-D. This led to modeling work by Liu et al. in Ref. [liu2019nf] and to a publication by Paz-Soldan et al. [pazsoldan2021nf], which included further MHD analysis of the DIII-D data as well as a comparison with the JET data published by Reux et al. In addition, Sheikh et al. contributed a study on the TCV and AUG tokamaks [sheikh2024ppcf], focusing on the operational windows for benign termination and the observed phenomenology with a brief discussion of measured MHD growth rates. In all these experiments, it was observed that there are limits to the amount of injected material for which the injection recombines the background, or companion, plasma. While on DIII-D this recombination was observed as a key component in making the mitigation both successful and benign, the corresponding density measurements on JET are not available for a number of discharges due to compromised diagnostic data and insufficient spatial resolution of the lines of sight of the interferometry systems. In addition, the interpretation of the pressure gauge signals was found to be hampered in these discharges [Schoonheere_2024]. However, all available data suggest that, also on JET, a successful recombination of the background plasma is the key requirement for a benign termination [reux2024runaway].
@@
First publications on the effect of material injections into RE beams in JET by Reux et al. used data collected before 2022 and, among other things, investigated the termination and mitigation characteristics in relation to the secondary injection material employed [reux2021prl, reux2022ppcf]. Not all of these early experiments were aimed at achieving a benign termination as defined by Paz-Soldan et al [pazsoldan2019ppcf]. In general, it was found that high-$Z$ or mixed secondary injections predominantly led to unsuccessful RE terminations. For such terminations, particularly with impurity fractions above 5%, a rich phenomenology was observed due to the variety of kinetic processes involved. Paz-Soldan et al. revisited the JET dataset and compared it with similar data from DIII-D [pazsoldan2021nf]. That work placed greater emphasis on investigating the MHD and equilibrium properties, although discharges with various secondary injection mixtures were also presented. At the time, it was concluded that successful and safe benign terminations were more likely to occur in the absence of high-$Z$ material. This finding motivated further JET experiments focusing on clean hydrogenic injections. The present publication differs from the work in Refs. [pazsoldan2021nf, reux2022ppcf] by restricting itself to the subset of JET and DIII-D discharges with hydrogenic injections and impurity fractions below 5%, which have a higher fraction of successful, benign terminations. However, low impurity fractions are not a sufficient criteria for a benign termination characteristic.
@@
Furthermore, all the JET results published so far in Refs. [pazsoldan2021nf, reux2022ppcf] were limited to discharges with pre-disruptive plasma currents below $1.5$ MA (JET discharge numbers below $100,000$). Recent experiments on JET in 2022 and 2023 have reached higher pre-disruptive plasma and RE currents than on other machines, up to $I_{p}\approx 3$ MA. They have revealed challenges in achieving benign termination at high pre-disruptive current values of $I_{p}\geq 2.5$ MA. This is illustrated in Fig. 1(a), which shows the pre-disruptive plasma current $I_{p}$ on the x-axis and the RE current prior to termination on the y-axis, with only non-benign, red cases for $I_{p}\approx 3.0$ MA on JET (r.h.s.). All cases with $I_{p}>1.5$ MA stem from JET campaigns in 2022 and 2023 and have not yet been published in peer-reviewed literature, though a short description can be found in Ref. [reux2024runaway]. In general, understanding the distinguishing features of these high-$I_{p}$ experiments on JET is critical for reactor-relevant extrapolation. In particular, studying the dynamics of the terminating MHD activity is crucial, as a conclusive explanation for what differentiates the nature of the terminating MHD event between benign and non-benign cases is still lacking. In addition, the referenced publications on DIII-D and JET, in particular by Paz-Soldan et al. [pazsoldan2021nf], highlighted a phenomenological difference with benign termination on DIII-D at the lowest edge safety factors (reproduced on AUG and TCV [sheikh2024ppcf]), while on JET, benign terminations at higher edge safety factors were also observed. This poses an open question of how observations from JET compare to those on DIII-D, TCV, and AUG, and how to unify the observed phenomenology.
@@ 2 Scenario and Methodology
Non-benign termination on JET #105794
Figure 3: Typical non-benign termination on JET (#105794), marking the primary disruption (0), the hydrogenic injection (1), the RE shorter plateau phase with plasma compression towards the center post (2), and the onset of the termination (3), without significant magnetic activity.
@@
Both approaches of driving the beam unstable can trigger a large-scale MHD event and terminate the RE beam either without (benign) or with (non-benign) localized impact of the RE beam on the first wall or divertor structure. This is observed after the onset of the termination (3), characterized by a drop in plasma current and safety factor, together with a spike in magnetic activity. In this work, most of the data points, if not otherwise explicitly mentioned, are sampled just prior to the termination. Specifically, this refers to the time point of the last equilibrium calculation before the magnetic signature rises, marking the onset of the terminating MHD event. On JET, the classification of benign and non-benign terminations is based on IR camera observations and the corresponding heat fluxes, as well as on the scheme developed by Loarte et al. [Loarte_2011], which tracks the conversion of magnetic energy into kinetic energy, a process that is more efficient in non-benign cases. Both is documented in Reux et al. [reux2022ppcf]. On DIII-D, due to the lack of comparable IR camera coverage, measured hard X-ray (HXR) spikes serve as a indicator distinguishing clean, benign terminations from non-benign terminations in which the RE beam strikes the wall, emitting HXR, as already applied in [pazsoldan2021nf]. It seems important to highlight that the community has not yet settled on a unified classification scheme for the characteristics of terminations and that, for consistency with previous work, this study adopts the respective definitions used on each machine. However, Paz-Soldan et al. verified the agreement between the HXR and IR classification schemes for selected experiments on DIII-D, as discussed in more detail in [pazsoldan2021nf]. A more unified classification scheme should be pursued in the future, but this is beyond the scope of this work.
A typical non-benign termination is shown in Fig. 3. As in the benign counterpart, Panel (a) shows the evolution of the plasma current. The entire RE phase is significantly shorter than in the benign case. The final loss event, see (3), is not abrupt but rather a slow decay of the REs, which are not efficiently deconfined. This can be seen in Panel (b), in which the magnetic activity during the termination is insignificant, indicating an inefficient MHD event in the non-benign case. The line-integrated density in Panel (c) suggests that the recombination of the free electrons was not successful, the material injection leads to a large increase of the measurement after (1). The signal drops steeply to zero at the termination when the residual plasma is lost and deconfined. The edge safety factor shown in Panel (d) illustrates the rapid compression of the RE beam and indicates that a low edge safety factor is reached prior to the onset of the termination at (3).
@@
A more detailed investigation of the MHD events in this work is based on fast magnetic measurements for MHD mode decomposition and equilibrium reconstruction. The primary diagnostics for the magnetic measurements are the Mirnov coils, which are small pickup coils placed around the tokamak vessel to measure time-varying magnetic fields. They operate on the principle of electromagnetic induction: a changing magnetic flux through the coil induces a voltage proportional to $\dot{B}$, the time derivative of the perturbed magnetic field. An example of such a measurement during the terminating MHD event of the JET discharge #102617 is shown in Panel (a) of Fig. 4 (solid line). Arrays of Mirnov coils distributed toroidally enable the mode decomposition of the measured perturbations in terms of the relative phases and amplitudes for different mode numbers $n$ through matrix operations [Munaretto_2021]. While DIII-D provides a large number of magnetic arrays, at JET only 5577 fast Mirnov coils were usable, depending on the plasma discharge. This restricts the analysis to low mode numbers $n=1,2$. Because of their fast response and sampling rates on the MHz scale, Mirnov coils are well suited for studying MHD instabilities, mode growth, and termination dynamics, where rapid changes in the magnetic field occur. Integration of the coil signals over the MHD event correlated to the final loss event yields $\delta B$, providing direct information on the amplitude of the underlying magnetic perturbations, as indicated by the shaded region under the curve in Panel (a). By fitting a linear curve, dotted line in Panel (b), to the logarithmic $\dot{B}$ values (solid line), the growth rate $\gamma$ can be measured. Fitting $\gamma$ with this approach is valid only during the initial exponential growth phase, when nonlinear coupling and saturation are not yet occurring. In this regime, fitting $\dot{B}$ or $\delta B$ yields identical results, differing only by constant offsets in the linear fit. The same constant offsets are introduced when considering conducting and shielding wall structures in front of magnetic sensors and, therefore, do not affect the results presented here. It should be noted that the Mirnov coil arrays used in this work measure the perturbed poloidal magnetic field $\delta B_{p}$, which, for simplicity, is referred to as $\delta B$ throughout this work. During the course of this work, the use of a normalization to the distance between the measurement array and the last closed magnetic flux surface was attempted, and the resulting $\delta B$ values were carefully compared. The normalized and un-normalized $\delta B$ values were found to exhibit an approximate 1:1 relationship, with some unsystematic scatter. The distance used for normalization averaged $1.04$ m (with a standard deviation of $0.35$ m) and was therefore not found to significantly alter the trends observed in the subsequent analysis. A consistent normalization would require knowledge of the RE beam position during the termination events, as well as information about the poloidal mode structure [igochine2015active], neither of which can be obtained in the studied scenario. Therefore, the authors refrain from applying any normalization. As a final caveat, it should be noted that significant efforts were made during the course of this project to automate the identification of the time window for the integration of $\delta B$ from the $\dot{B}$ signal, as well as the fitting window for the exponential growth rates. However, due to the large variety of mode characteristics encountered, such as non-linear damping, mode rotation, and $n=0$ pollution, a very careful manual identification of the integration and fitting time windows was ultimately required. This process relied on identifying the first strong $\dot{B}$ peak correlated with RE deconfinement and integrating from the associated rise of the left flank out of the noise to its decline. Establishing automated workflows for large database studies would require dedicated analysis projects and is beyond the scope of this work.
@@ 3.1 Evolution of high-$I_{p}$ RE beams on JET
The first critical observation concerns the fundamental difference in the duration of the RE plateau phase between benign and non-benign cases. In typical benign scenarios, the RE plateau phase is long, controlled, and stable, with an average duration of $0.62$ s (standard deviation $0.66$ s). In contrast, in cases where recombination is unsuccessful and the outcome is non-benign, no stable plateau forms, with an average duration of $0.07$ s (standard deviation $0.03$ s) before the termination. This is likely due to the higher resistivity of the non-recombined plasma, which results in a situation where the spatial position of the RE beam is no longer effectively controlled. The subsequent MHD response, however, is assumed to be largely indifferent to the specific pathway by which the system becomes unstable. In addition, typical MHD time scales, as discussed later, are on the order of microseconds, such that this deviation in RE plateau length cannot systematically interact with the time the MHD instabilities have to develop. A more detailed investigation of how the RE trajectory influences key quantities, such as the current profile distribution, and, consequently, the MHD instability is left for future work and cannot be disentangled from present experiments.
@@
The next step of the experimental analysis concerns the properties of the plasma equilibrium. In Fig. 5, the trajectory of the reconstructed magnetic axis is plotted for the last 300 ms before termination in the $R$-$z$-space with the JET contour in black for comparison. For cases in which the RE plateau phase was shorter, sampling was performed from the onset of the RE plateau phase. The panel on the left shows the non-benign cases, while the panel on the right shows the benign cases. The color coding corresponds to the cross-section of the RE column. It is observed that the onset of termination in non-benign cases occurs at a smaller cross-section and in a less central position, whereas benign terminations typically begin from a relatively larger cross-section and a more central position. This suggests a different evolution of the RE beam in benign and non-benign cases, with non-benign beams being compressed to smaller cross-sections and, consequently, to lower edge safety factors than their benign counterparts. As already noted, the non-benign cases on JET are usually very short, suggesting that the higher resistivity of non-recombined plasmas likely affects their controllability. Under these circumstances, the short lifetime of the non-recombined plasmas could interact with the evolution of the current profiles. Assuming identical current-profile evolution across shots, for example through competing time scales of different RE generation or current diffusion and redistribution mechanisms, any systematic variation in duration would alter the current profile at the point of the secondary disruption. A realistic modeling of the RE evolution and formation, as already stated, is the subject of a subsequent publication on this topic.
@@
As discussed above, the non-benign cases on JET are associated with high pre-disruptive plasma currents. The higher stored energy in these cases leads to strong RE currents. Combined with the observed smaller cross-sections, this results in higher RE current densities prior to termination. Consequently, a clear separation between benign and non-benign cases emerges in terms of their RE current density $j_{RE}=I_{RE}/A$, as illustrated in Fig. 1(b) with the JET datapoints on the r.h.s. This behavior is not commonly observed across the previously discussed devices, and a similar analysis for the studied DIII-D dataset does not show such a separation, see l.h.s. of the same figure. This suggests that different mechanisms play a role in the formation or termination of the RE currents on the two machines. Interestingly, at the highest $I_{p}\approx 1.2$ MA on DIII-D, the only non-benign case does separate from the benign cases with thresholds similar to those in the JET dataset, but more data would be needed to draw firm conclusions. For JET, however, the higher RE current densities may interact with kinetic effects, as suggested by Hollmann and Hoppe, who noted that more concentrated (peaked) RE current density profiles may hinder effective recombination of the companion plasma or can even lead to re-ionization of an initially recombined background plasma [Hollmann2023, Hoppe2025].
@@ 3.2 Intermittent MHD events prior to the termination in JET
In Fig. 7(a), the measured perturbed $\delta B$ is shown for benign and non-benign cases that pass through or terminate at $q_{\text{edge}}=3$. The MHD amplitude in benign cases increases linearly with the RE current deconfined during the event. This suggests that there exists a threshold in $\delta B/B$ at which the MHD mode becomes strong enough to trigger a catastrophic, terminating MHD event. Here, $B$ represents the poloidal magnetic field, with $B_{p}\sim I_{RE}$ since $\delta B=\delta B_{p}$ and the variation in the toroidal magnetic field $B_{t}$ is small in this dataset, ranging only from $3.0$ to $3.45$ T. Non-benign cases exhibit a clearly suppressed MHD amplitude, insufficient to cause a proper termination. At the moment, it remains an open question how the recombination state of the companion plasma influences the amplitude or saturation of these MHD events.
A possible explanation for the observed suppression is provided in Panel (b), which shows a tendency toward higher $l_{i}$ values in most of the non-benign cases, suggesting a more peaked RE current distribution. The intermittent small events do not change the measured $l_{i}$ values in the final phase prior to termination significantly. Variations in the current profiles could originate from the dependence of the RE current density distribution on the specific mechanisms by which the RE current forms during the primary disruption. JET, in particular, is characterized by strong RE avalanche generation, which could explain the different $j_{RE}$ profiles observed. This was discussed in detail by Smith et al. [Smith_2010] who predicted that the absolute pre-disruptive current can influence the RE current distribution on JET. Steepening of the RE current profiles, correlated with higher pre-disruptive $I_{p}$, was also observed on JET by Loarte et al. [Loarte_2011].